Solve.
d²y/dx² - 4x dy/dx + (4x²-1)y = -3exp(x²)sin(2x)
1
Expert's answer
2019-08-25T14:47:19-0400

"y^{\prime\prime}-4xy^{\prime}+(4x^2-1)y=-3e^{x²}sin(2x)\\" consider:"y^{\prime\prime}-4xy^{\prime}+(4x^2-1)y=0\\
y(x)=e^{x^2}\cdot z(x)\\
(2e^{x^2}z+4x^2e^{x^2}z+4xe^{x^2}z^{\prime}+e^{x^2}z^{\prime\prime})-4x(2xe^{x^2}z+e^{x^2}z^{\prime})+(4x^2-1)e^{x^2}z=0\\
z^{\prime\prime}+z=0\\
z=c_1\cos(x)+c_2\sin(x)\\
y(x)=c_1e^{x^2}\cos(x)+c_2e^{x^2}\sin(x)"

check the particular solution:"y_1=e^{x^2}\sin(2x)\\
y^\prime = 2\,x{{\rm e}^{{x}^{2}}}\sin \left( 2\,x \right) +2\,{{\rm e}^{{x}^{2}}
}\cos \left( 2\,x \right)\\
y^{\prime\prime} = -2\,{{\rm e}^{{x}^{2}}}\sin \left( 2\,x \right) +4\,{x}^{2}{{\rm e}^{{
x}^{2}}}\sin \left( 2\,x \right) +8\,x{{\rm e}^{{x}^{2}}}\cos \left( 2
\,x \right) \\
y^{\prime\prime}-4xy^{\prime}+(4x^2-1)y =\\
-2\,{{\rm e}^{{x}^{2}}}\sin \left( 2\,x \right) +4\,{x}^{2}{{\rm e}^{{
x}^{2}}}\sin \left( 2\,x \right) +8\,x{{\rm e}^{{x}^{2}}}\cos \left( 2
\,x \right) -4x(2\,x{{\rm e}^{{x}^{2}}}\sin \left( 2\,x \right) +2\,{{\rm e}^{{x}^{2}}
}\cos \left( 2\,x \right) )+(4x^2-1)(e^{x^2}\sin(2x))=-3e^{x^2}\sin(2x)"

then solution is:"y(x)=c_1e^{x^2}\cos(x)+c_2e^{x^2}\sin(x)+e^{x^2}\sin(2x)"